OpenStudy (anonymous):

find the equation of the lines making an angle of tan^-1(3) with the line 2x-y-8=0 and passing through (4,-1)

OpenStudy (cwrw238):

y = 2x - 8 has slope 2 nad lakes an angle of 63.4 degrees with horizontal the other line will make angle tan-1(3) = 71.6 to this line that is 135 degrees to horizontal which is a slope of -1 (tan 135 = -1) it also goes through the point (4,-1) so using the standard form y-y=m(x-x1) we have y-(-1) = -1(x - 4) y + 1 = -x + 4 y + x = 3

OpenStudy (anonymous):

Thank you for your help, it is very much appreciated.